# Simulating non-Brownian suspensions with non-homogeneous Navier slip boundary conditions

### v1

@@ Abstract
Fluid-structure interactions are commonly modeled using no-slip boundary conditions. However, small deviations from these conditions can significantly alter the dynamics of suspensions and particles, especially at the micro and nano scales. This work presents a robust implicit numerical method for simulating non-colloidal suspensions with non-homogeneous Navier slip boundary conditions. Our approach is based on a regularized boundary integral formulation, enabling accurate and efficient computation of hydrodynamic interactions. This makes the method well-suited for large-scale simulations. We validate the method by comparing computed drag forces on homogeneous and Janus particles with analytical results. Additionally, we consider the effective viscosity of suspensions with varying slip lengths, benchmarking against available analytical no-slip and partial-slip theories.
@@ 1 Introduction
In addition to experimental and theoretical studies, numerical methods have been used to investigate systems with partial slip BCs. These methods can be broadly categorized into explicit and implicit methods. Explicit methods directly resolve the fluid flow and include molecular dynamics (MD) [16, 23, 24], smoothed particle hydrodynamics (SPH) [25], smoothed dissipative particle dynamics (SDPD) [26], and dissipative particle dynamics (DPD) [27], among others. While effective, these methods are computationally demanding due to their need to discretize the fluid explicitly. In contrast, implicit methods do not solve for the flow field directly. Instead, they use the Green’s functions of the Stokes equations to compute the velocities of solute particles. These methods are limited to the study of suspensions in the Stokes regime. However, with this methods only the particles surfaces are discretized which reduces memory requirements considerably. Some implicit analytical methodologies have been also developed to include partial slip BCs [28, 29, 30, 9, 22].
This work presents a robust implicit method for modeling particles with partial slip BCs, capable of simulating complex and long-range interactions in particle suspensions. Additionally, the particles can feature non-homogeneous slip lengths along their surfaces. Our method employs a regularized boundary integral formulation to implement partial slip BCs and compute hydrodynamic tractions and particle velocities [31]. The approach achieves high accuracy, with only a few percent relative errors, even with a modest discretization per particle. Thus, it is appropriate for simulating large-scale particulate systems.
@@ 2 Continuum formulation
where $\boldsymbol{v}$ and $p$ are the flow velocity and pressure and $\eta$ the fluid viscosity. The BCs on the particles surfaces are model as a slip condition
@@
We make use of an integral formulation to solve the Stokes equations [35]. First, we introduce the the single layer operator, $\boldsymbol{\mathcal{M}}$, acting on the hydrodynamic traction $\boldsymbol{\lambda}$, and the Stokes double layer operator, $\boldsymbol{\mathcal{D}}$, acting on the fluid velocity
@@
where $\boldsymbol{\mathcal{I}}$ is the identity operator. Replacing the slip condition, Eq. 3, one obtains [23, 33]
@@ 3 Discretization
Combining the four equations (12)-(15) gives the linear system
@@ 3.1 Numerical solver
Figure 2: Convergence of preconditioned GMRES to solve the mobility problem for a single particle discretized with a different number of blobs. The figure shows results for no-slip (left) and full-slip (right) BCs.
Figure 3: GMRES convergence for $\boldsymbol{M}$ particles, from 1 to 4096, forming a simple cubic lattice with a distance between first neighbors equal to two particle diameters. The particles are discretized with $N_{b}$ = 42 blobs.
The linear system (17) can be large for many-particle suspensions, therefore, using a direct solver is computationally unfeasible. Instead, we solve it with an iterative method. Since (17) is non-symmetric we employ a preconditioned GMRES solver which has been shown to be a very efficient for similar Stokes problems [38, 39].
To converge within a moderate number of iterations it is necessary to employ a preconditioner which approximates the inverse of (17). Since the preconditioner speeds up the convergence but it does not modify the solution of the linear system we can make a number of approximation so it can be applied efficiently. First, we note that in the continuum limit the action of the double layer on rigid body velocities is equivalent to applying half the identity operator [35]. Thus, on the preconditioner we approximate $\left(\frac{1}{2}\boldsymbol{I}+\boldsymbol{D}\right)\boldsymbol{K}\approx\boldsymbol{K}$. Second, in the mobility matrix we neglect the interaction between blobs belonging to different bodies, i.e. we replace the mobility matrix $\boldsymbol{M}$ by its block-diagonal approximation $\widetilde{\boldsymbol{M}}^{(pq)}=\delta_{pq}\boldsymbol{M}^{(pp)}$ where the superscripts $p$ and $q$ indicate the block of blobs belonging to body $p$ or $q$ respectively. Finally, we simply neglect the action of the double layer on the slip velocity. Applying the preconditioner derived with these approximations is equivalent to solve the system
@@
We show in Fig. 2 the convergence of GMRES to solve the mobility problem for a single particle discretized with different number of blobs. We see that regardless of the discretization and the slip length the solver converges quickly, within a small number of iterations. In Fig. 3 we show the convergence for particles, discretized with $42$ blobs, forming a simple cubic lattice with distances betweeen first neighbors of two particle diameters. The figure shows that the convergence depends very weakly on the number of particles even for very large systems.
@@ 4.3 Janus particles
Figure 5: Janus particles have one half (colored yellow) representing a no-slip surface, while the other half (colored purple) represents a full-slip surface. The interphase line region is an average slip length. The scheme illustrates examples of Janus particles aligned along the P. axis at three angles: 0, 45, and 90 degrees from the horizontal x-axis. In all cases, the applied force acts along the horizontal x-axis.
@@ 4.4 Suspension viscosity
In the continuum limit, the last two terms in (25) cancel out, thus the continuum problem is not modified. In the discrete case, for constant background flow, the first two terms cancel exactly and the last term balances the term $\left(\frac{1}{2}\boldsymbol{I}+\boldsymbol{D}\right)\boldsymbol{K}\boldsymbol{U}$ in (24) which guarantees that particles are advected with the background velocity to machine precision. Using this background flow it is possible to obtain accurate viscosity results as we will show in a moment.
The second detail to extract the viscosity with higher accuracy is to measure the flow velocity correctly. Since we solve the Stokes equations with an implicit method, we do not have direct access to the flow (although it can be computed). Therefore, we measure the average particle velocities profile, $\langle u_{x}\rangle(z)$, and use the Faxén law to extract the ambient velocity magnitude from
@@
Finally, in the right panel of Fig. 7 we show the suspension viscosity versus volume fraction for several slip lengths. The no-slip results agree well with the Stokesian dynamics results of Sierou and Brady [34], while the viscosity grows slower for larger slip lengths as expected. The viscosity decreases with the slip length until $\ell/R\approx 10$. Further increases in the slip length do not reduce the viscosity any lower. We also compare our numerical results with an approximated theory derived from a Stokes–Darcy model with partial slip [12]. For small slip lengths (e.g., $\ell\leq 0.1$), the numerical results show good agreement with the analytical predictions. However, for larger slip lengths, the theory proposed in Ref. [12] consistently overpredicts the effective viscosity. This overprediction occurs across all volume fractions, as the theory’s prediction for low concentrations of free-slip particles, $\eta_{\text{eff}}/\eta=1+1.6\phi$, exceeds the exact prediction, to linear order, given by Eq. (27). Nevertheless, the theory successfully captures the overall trend observed in the numerical results.
@@ 5 Conclusions
We have validated the accuracy of our scheme across a range of slip lengths, from full-slip to no-slip BCs. In all cases, our results remained accurate within a few percent, even with resolutions as low as 42 blobs per particle. Additionally, we have reproduced the drag forces and torques on Janus particles. We have also measured the high-frequency viscosity in suspensions of particles with homogeneous slip distributions. At low volume fractions, our viscosity results are in good agreement with analytical predictions [2]. Over a broad range of volume fractions, our results are consistent with Stokesian dynamics simulations in the limit of small slip lengths ($\ell/R\approx 0.01$) while, as the slip length increases, the viscosity decreases as expected. We found that the viscosity decreases with the slip length up to $\ell/R=10$, beyond which it becomes independent of the slip length.
The scheme presented in this work is flexible enough to model inhomogeneous slip length distributions, incorporating for instance, regular and irregular patterns over the surfaces, that have potential applications in nanotechnology and biomedicine [43, 44, 45]. The extension of this work to Brownian suspensions would be an important contribution. However, how to generate Brownian velocities when the double layer operator, which is non-symmetric and thus not positive definite, appears in Eq. (11) remains an open problem. Our implementation of the method is publicly available as free software at https://github.com/stochasticHydroTools/RigidMultiblobsWall.

### v2

@@ Abstract
Fluid-structure interactions are commonly modeled using no-slip boundary conditions. However, small deviations from these conditions can significantly alter the dynamics of suspensions and particles, especially at the micro and nano scales. This work presents a robust implicit solvent method for simulating non-colloidal suspensions with non-homogeneous Navier slip boundary conditions. Our approach is based on a regularized boundary integral formulation, enabling accurate and efficient computation of hydrodynamic interactions. This makes the method well-suited for large-scale simulations. We validate the method by comparing computed drag forces on homogeneous and Janus particles with analytical results. Additionally, we consider the effective viscosity of suspensions with varying slip lengths, benchmarking against available analytical no-slip and partial-slip theories.
@@ 1 Introduction
In addition to experimental and theoretical studies, numerical methods have been used to investigate systems with partial slip BCs. These methods can be broadly categorized into explicit and implicit solvent methods. Explicit solvent methods directly resolve the fluid flow and include molecular dynamics (MD) [16, 23, 24], smoothed particle hydrodynamics (SPH) [25], smoothed dissipative particle dynamics (SDPD) [26], and dissipative particle dynamics (DPD) [27], immersed boundary method (IBM) [28], among others. The explicit discretization of the fluid makes these methods versatile, as they can deal with, for example, non-constant viscosity [29, 30] but also computationally demanding. In contrast, implicit solvent methods do not solve for the flow field directly. Instead, they use the Green’s functions of the Stokes equations to compute the velocities of solute particles. These methods are limited to the study of suspensions in the Stokes regime. However, with this methods only the particles surfaces are discretized which reduces memory requirements considerably. Some implicit analytical methodologies have been also developed to include partial slip BCs [31, 32, 33, 9, 22].
This work presents a robust implicit solvent method for modeling particles with partial slip BCs, capable of simulating complex and long-range interactions in particle suspensions. Additionally, the particles can feature non-homogeneous slip lengths along their surfaces. Our method employs a regularized boundary integral formulation to implement partial slip BCs and compute hydrodynamic tractions and particle velocities [34]. The approach achieves high accuracy, with only a few percent relative errors, even with a modest discretization per particle. Thus, it is appropriate for simulating large-scale particulate systems.
@@ 2 Continuum formulation
where $\boldsymbol{v}$ and $p$ are the flow velocity and pressure and $\eta$ the constant fluid viscosity. The BCs on the particles surfaces are model as a slip condition
@@
[…]
@@
[…]
@@ 3 Discretization
The four equations (12)-(15) form a linear system for the particles velocities, tractions and slip velocities. This linear system can be written in compact notation with the help of the geometric matrix $\boldsymbol{K}$, which transform rigid body velocities into surface velocities
$\boldsymbol{K}\begin{bmatrix}\boldsymbol{u}_{m}\\ \boldsymbol{\omega}_{m}\end{bmatrix}=\boldsymbol{u}_{m}+\boldsymbol{\omega}_{m}\times(\boldsymbol{r}_{j}-\boldsymbol{q}_{m})\text{ for }j\in\mathcal{B}_{m}.$ (17)
Then, combining the four equations (12)-(15) and using (17) gives the linear system
@@ 3.1 Numerical solver
Figure 2: Convergence of preconditioned GMRES to solve the mobility problem Eq. (18) for a single particle discretized with a different number of blobs. The figure shows results for no-slip (left) and full-slip (right) BCs.
Figure 3: GMRES convergence to solve the mobility problem Eq. (18) for $\boldsymbol{M}$ particles, from 1 to 4096, forming a simple cubic lattice with a distance between first neighbors equal to two particle diameters. The particles are discretized with $N_{b}$ = 42 blobs.
The linear system Eq. (18) can be large for many-particle suspensions, therefore, using a direct solver is computationally unfeasible. Instead, we solve it with an iterative method. Since Eq. (18) is non-symmetric we employ a preconditioned GMRES solver which has been shown to be a very efficient for similar Stokes problems [41, 42].
To converge within a moderate number of iterations it is necessary to employ a preconditioner which approximates the inverse of Eq. (18). Since the preconditioner speeds up the convergence but it does not modify the solution of the linear system we can make a number of approximation so it can be applied efficiently. First, we note that in the continuum limit the action of the double layer on rigid body velocities is equivalent to applying half the identity operator [38]. Thus, on the preconditioner we approximate $\left(\frac{1}{2}\boldsymbol{I}+\boldsymbol{D}\right)\boldsymbol{K}\approx\boldsymbol{K}$. Second, in the mobility matrix we neglect the interaction between blobs belonging to different bodies, i.e. we replace the mobility matrix $\boldsymbol{M}$ by its block-diagonal approximation $\widetilde{\boldsymbol{M}}^{(pq)}=\delta_{pq}\boldsymbol{M}^{(pp)}$ where the superscripts $p$ and $q$ indicate the block of blobs belonging to body $p$ or $q$ respectively. Finally, we simply neglect the action of the double layer on the slip velocity. Applying the preconditioner derived with these approximations is equivalent to solve the system
@@
We show in Fig. 2 the convergence of GMRES to solve the mobility problem Eq. (18) for a single particle discretized with different number of blobs. We see that regardless of the discretization and the slip length the solver converges quickly, within a small number of iterations. In Fig. 3 we show the convergence for particles, discretized with $42$ blobs, forming a simple cubic lattice with distances between first neighbors of two particle diameters. The figure shows that the convergence depends very weakly on the number of particles even for very large systems.
@@ 4.3 Janus particles
Figure 5: Janus particles have one half (colored yellow) representing a no-slip surface, while the other half (colored purple) represents a full-slip surface. The interphase line region is an average slip length. The scheme illustrates examples of Janus particles aligned along the particle axis at three angles: 0, 45, and 90 degrees from the horizontal x-axis. In all cases, the applied force acts along the horizontal x-axis.
@@
4.4 Random slip distribution
The current method allows for more complex slip distributions. To showcase the flexibility of the method, we measure the translational and rotational drag on spheres with random slip distributions. We take a spherical particle discretized with $N_{b}=42$ blobs and set the slip on the blobs to either full-slip ($\ell/R=10^{3}$) or no-slip ($\ell/R=10^{-3}$), varying the full-slip coverage from 0% to 100%. For each slip coverage we use 10 random slip distributions and measure the average drag and the standard deviation, see Fig. 7bc. Interestingly, the results show a nonlinear variation between the full no-slip and full-slip distributions, while the drag standard deviations, shown as error bars in Fig. 7bc, are small. The translational drag standard deviation is below 5% for all slip coverages. For 50% full-slip coverage spheres with random slip distributions have a drag about 10% larger than a Janus sphere, whose drag is shown as a red star in Fig. 7b.
Next, we study the convergence of the numerical results with increased resolution. For each slip arrangement, we increase the resolution to 162 and 642 blobs while maintaining the same slip distribution, see Fig. 7a. The drag computed with the refined resolutions agrees well with the coarse resolution. Thus, coarse resolutions can be used as long as the blob radius is similar to, or smaller, than the characteristic length of the spatial slip distribution.
Figure 7: (a) Example of a sphere discretized with $N_{b}=42$ blobs and a random slip distribution and refined discretizations with 162 and 642 blobs and the same slip distribution. (b,c) Translational and rotational drag versus slip coverage from full no-slip to full-slip computed for three resolutions. For each coverage 10 random slip distributions are used and the standard deviation in the drag is shown as error bars. The red star in (b) represents the drag of a Janus sphere discretized with $N_{b}=10242$ blobs as in Sec. 4.3.
4.5 Spherocylinders
In this section, we study the capability of our numerical method to model spherocylindrical particles. We use spherocylinders formed by a cylindrical body of length $L$ and radius $R=L/2$, with two spherical caps of radius $R$ at the ends. First, we consider Janus particles with an axisymmetric slip length distribution. The surface of a Janus particle is divided into two domains: one with no-slip ($\ell/L=10^{-3}$) and the other with full-slip ($\ell/L=10^{3}$). The position of the interface between the two domains is set by the polar angle $\alpha$, see Fig. 8a. We measure the translational and rotational drag as $\alpha$ varies smoothly from $\alpha=0$ to $180^{\circ}$, i.e., from a particle with homogeneous no-slip boundary conditions to a particle with homogeneous full-slip boundary conditions. To calculate the translational drag, we measure the force on a particle moving with unit velocity along, or perpendicular, to its main axis. Similarly, we measure the rotational drag when it rotates with unit angular velocity around, or perpendicular, to its main axis.
Figure 8: (a) Spherocylinders of aspect radio 44 discretized (from left to right) with $N_{b}=132,\,583,\,2364,\text{ and }9929$ blobs. The polar angle $\alpha$ sets the interface between the domain with full-slip (dark blobs) or no-slip (light blobs). (b) Spherocylinder’ translational drag parallel to its main axis (continuous lines) and perpendicular (dashed lines) versus the $\alpha$. (c) Rotational drag, with the same notation as in (b).
We present the drag results obtained with resolutions between $N_{b}=132$ and $N_{b}=9929$ blobs in Fig. 8bc. All the drag components decay with $\alpha$, i.e., with increasing full-slip coverage. The largest change appears in the rotational drag around the main axis. In that case, the drag decays to zero for a homogeneous full-slip rod because all the fluid motion is tangential to the surface. In the other cases, the flow has non-tangential components, and thus the drag never decays to zero, even for the homogeneous full-slip case. The results show that a coarse resolution of $N_{b}=132$ blobs is sufficient to obtain results accurate within 10% compared to the refined results for all cases except for the homogeneous full-slip case, where the errors are around 15% for the translational drag along the main axis. Interestingly, the rotational drags have small errors for all slip distributions.
Figure 9: (a) Sketch of a spherocylinder in a shear flow and definition of $\theta$, the orientation angle with the flow. (b) Angular velocity of a spherocylinder in a shear flow respect to the angle $\theta$ formed by the cylinder main axis and the direction of the flow. For small slip lengths the angular velocity is always positive while for large slip lengths the angular velocities become negative for particles nearly aligned with the shear flow. (c) Equilibrium angle with the shear flow respect the slip length.
As a second test, we study how the angular velocity of a spherocylinder with the same geometry ($R=L/2$) and a homogeneous slip distribution, immersed in a shear flow $\boldsymbol{v}_{0}=\dot{\gamma}z\boldsymbol{e}_{x}$, varies with the slip length. Since we showed in the previous test that the rotational drag was recovered with good accuracy using a coarse discretization, in this test we use a particle discretized with $N_{b}=132$ blobs, see Fig. 9a. We show in Fig. 9b the angular velocity of the particle with respect to its orientation with the flow for several slip lengths. For small slip lengths, the angular velocity is always positive, i.e., the particle rotates with the flow as expected. Interestingly, for slip lengths larger than $\ell/L\approx 1/2$, the angular velocity becomes negative for some orientations, i.e., the particle counter-rotates with respect to the flow. The sign flip in the angular velocity indicates that particles with sufficiently large slip lengths can reach an equilibrium orientation with respect to the shear flow. We show in Fig. 9c the equilibrium angle for several slip lengths. The equilibrium angle increases monotonically with $\ell$ and reaches a plateau for $\ell/L\approx 100$. This curious phenomenon, where a particle aligns in a shear flow, has been observed in graphene nanoplatelets, which exhibit very large slip lengths [23, 36]. We show here that the same effect appears for rod-like particles with sufficiently large slip lengths.
@@ 4.6 Suspension viscosity
Figure 10: Error in the angular velocity of a single spherical particle immersed in a shear flow versus the number of blobs $N_{b}$ used to discretize the sphere with and without the background flow correction Eq. (26). With the correction the error is about three times smaller.
@@
In the continuum limit, the last two terms in Eq. (26) cancel out, thus the continuum problem is not modified. In the discrete case, for constant background flow, the first two terms cancel exactly and the last term balances the term $\left(\frac{1}{2}\boldsymbol{I}+\boldsymbol{D}\right)\boldsymbol{K}\boldsymbol{U}$ in Eq. (25) which guarantees that particles are advected with the background velocity to machine precision. This correction also reduces the error in the angular velocity of a single spherical particle immersed in a shear flow as shown in Fig. 10. Using this background flow it is possible to obtain accurate viscosity results as we will show in a moment.
The second detail to extract the viscosity with higher accuracy is to measure the flow velocity correctly. Since we solve the Stokes equations with an implicit solvent method, we do not have direct access to the flow (although it can be computed). Therefore, we measure the average particle velocities profile, $\langle u_{x}\rangle(z)$, and use the Faxén law to extract the ambient velocity magnitude from
@@
Finally, in the right panel of Fig. 11 we show the suspension viscosity versus volume fraction for several slip lengths. The no-slip results agree well with the Stokesian dynamics results of Sierou and Brady [37], while the viscosity grows slower for larger slip lengths as expected. The viscosity decreases with the slip length until $\ell/R\approx 10$. Further increases in the slip length do not reduce the viscosity any lower. We also compare our numerical results with an approximated theory derived from a Stokes-Darcy model with partial slip [12]. For small slip lengths (e.g., $\ell\leq 0.1$), the numerical results show good agreement with the analytical predictions. However, for larger slip lengths, the theory proposed in Ref. [12] consistently overpredicts the effective viscosity. This overprediction occurs across all volume fractions, as the theory’s prediction for low concentrations of free-slip particles, $\eta_{\text{eff}}/\eta=1+1.6\phi$, exceeds the exact prediction, to linear order, given by Eq. (28). The theory from Ref. [12] derives the viscosity modeling the flow near a particle with a Stokes-Darcy model and perhaps the additional dissipation introduced by the porosity of the Darcy model overpredicts the viscosity. Nevertheless, the theory successfully captures the overall trend observed in the numerical results.
@@ 5 Conclusions
We have validated the accuracy of our scheme across a range of slip lengths, from full-slip to no-slip BCs. In all cases, our results remained accurate within a few percent, even with resolutions as low as 42 blobs per particle. Additionally, we have reproduced the drag forces and torques on Janus particles. We have also measured the high-frequency viscosity in suspensions of particles with homogeneous slip distributions. At low volume fractions, our viscosity results are in good agreement with analytical predictions [2]. Our results in the limit of small slip lengths ($\ell/R\approx 0.01$) are consistent with no-slip Stokesian Dynamics simulations over a broad range of volume fractions [37]. Then, as the slip length increases, the viscosity decreases as expected. We found that the viscosity decreases with the slip length up to $\ell/R=10$, beyond which it becomes independent of the slip length.
The use of an implicit solvent method reduces the memory requirements considerably as only the particles surfaces are discretized. For example, in the viscosity measurements of Sec. 4.6 we employed 250 particles each of then discretized with 162 blobs, thus, $40,500$ blobs in total. This number of degrees of freedom is independent of the volume fraction, $\phi$, as the solvent itself is not explicitly discretized. With those degrees of freedom one mobility problem was solved within 6 to 20 iterations to a tolerance $10^{-6}$ in about one minute per time step using only 8 cores per simulations. Thus, it is possible to run large simulations in a single workstation. The same system solved with an explicit solvent method would have required about $(L/(2a))^{3}$ particles to discretize the system, where $L$ is the system size and $a$ the blob radius or equivalently the discretization length scale. Using the same resolution as in our simulations ($R/a\approx 8$) that would be between $1.5\cdot 10^{5}$ and $6\cdot 10^{6}$ particles for volume fractions between $\phi=0.4$ and $\phi=0.01$, like the ones used in Fig. 11. Thus, low volume fraction simulations become particularly expensive in explicit solvent methods. On the other hand, inertial effects or non-Newtonian solvent properties require a full description of the solvent domain [29, 30].
The scheme presented in this work is flexible enough to model inhomogeneous slip length distributions, incorporating for instance, regular and irregular patterns over the surfaces, that have potential applications in nanotechnology and biomedicine [46, 47, 48]. The extension of this work to Brownian suspensions would be an important contribution. However, how to generate Brownian velocities when the double layer operator, which is non-symmetric and thus not positive definite, appears in Eq. (11) remains an open problem. For no-slip boundary conditions it is possible to generate Brownian displacements by including in the right-hand-side of the mobility problem (Eq. (18)) a random flow compatible with the Stokes equations [49]. The physical interpretation is that the particles diffuse by the motion generated by the random flow. Using the fact that the mobility matrix $\boldsymbol{M}$ is positive definite, it is possible to prove that this approach obeys the fluctuation-dissipation balance [49]. This suggest that the same idea should work for partial slip boundary conditions, however, the presence of double layer operator in the mobility problem breaks the fluctuation-dissipation balance. Bedeaux, Albano and Mazur studied the formulation of Brownian dynamics with non-zero slip lengths [50]. They showed that in the context of fluctuating hydrodynamics it is necessary to include an additional random frictional force at the fluid-particle interface to balance the dissipation generated by the slip and fulfill the fluctuation-dissipation balance [50]. Further exploration of this approach will be left for future work. Our implementation of the method is publicly available as free software at https://github.com/stochasticHydroTools/RigidMultiblobsWall.